for what value of x does this system has non-trivial solution,x-y+2z=0,2x+y+z=0,2x+y+2z=0
1
Expert's answer
2012-07-17T09:00:11-0400
y=-z-x then z=-(2/3)x y=-(4/3)x. This solution is not trivial, when x <>0.